OpenStudy (anonymous):

Find two positive numbers whose product is 400 and such that the sum of twice the first and three times the second is a minimum. Optimization Problem

OpenStudy (anonymous):

equations are x y = 400 minimize 2 x + 3 y convert problem to single variable using constraint y = 400/x rewrite goal minimize f(x) = 2 x + 3 (400/x) set derivative =0 f'(x) = 2 -1200 x^-2 [two minus 1200 divided by square of x] 2 x^2 = 1200 x = sqrt(600) = 24.5 y = 400/x= 16.3 min f = 2(24.5) + 3(16.3) = 97.9 to be sure this is a minimum, take second derivative and show that it is positive over the domain. f"(x) = 2400 x^-3 positive for x>0 Does check. That's the method. Try it yourself to be sure I have not calculated anything wrong.

OpenStudy (anonymous):

Thanks